OpenStudy (hexiflexidexi):

Vector help: Find an equation of the plane parallel to 2x + y − z = 1 and passing through the point (0, 2, −2). I don't need an answer, I just need you to guide me through the process--actually, don't give me the answer, just guide me on the way... Thanks!

OpenStudy (mathmale):

Planes are specified by their normal vectors in combination with specification of a point in the plane. Figure out how you could create a vector nomral to the plane. 2x + y -z = 1 is parallel to this unknown plane, meaning that 2x + y - z = 1 will never intersect the plane. Hint: you'll need to find the cross product of two vectors which are themselves in or parallel to the plane.

OpenStudy (mathmale):

Welcome to OpenStudy!

OpenStudy (hexiflexidexi):

Thanks :)

OpenStudy (hexiflexidexi):

@mathmale To find vectors on the plane/parallel to the plane, do I substitute numbers into x, y, and z?

OpenStudy (mathmale):

No. You are given the equation of a line parallel to the plane. From this equation you can find the x-, y- and z-components of a vector that is parallel to the plane. Now let (x, y, z) be a point lying in the plane. Find the vector that goes from this point to the given point in the plane. Take the cross product of these two vectors. Voila! You'll have the equation of the plane.

OpenStudy (hexiflexidexi):

@mathmale Uh, a few points I don't understand (sorry~) 1. "From this equation you can find the x-, y- and z-components of a vector that is parallel to the plane." So, like there is a vector lying on the give plane which is parallel to the plane that we want? Also, how can I derive a vector from a an equation of the plane? 2. "Now let (x, y, z) be a point lying in the plane. Find the vector that goes from this point to the given point in the plane." Does (x,y,z) lie on the plane that we want, or from the given plane? (Also, to make sure I got the facts right, the vector between the points can be calculated by x,y,z values of vector1-vector2?) Thank you for being so patient, I've just started learning this and can be a bit...dense. :)

OpenStudy (gottennis121):

@Keigh2015 @hakunamatata34 @SeanStewart16

OpenStudy (hakunamatata34):

yes

OpenStudy (hexiflexidexi):

Hey, I'm not sure what to do next for this <3

OpenStudy (hakunamatata34):

im sorry idk this one:(

OpenStudy (hexiflexidexi):

@mathmale

OpenStudy (mathmale):

2x + y − z = 1 is a given line. What is its direction vector? We are told that it is parallel to the plane. As I said before, this line will never intersect the plane. You could imagine moving this line up or down in a direction perpendicular to the plane until the line actually lies IN the plane. Next, you create a new vector, a vector that extends from a point (x, y, z) in the plane to the given point in the plane. One way of creating such a vector is represented by the following:\[<x-x _{1},y-y _{1},z-z _{1}>, \]

OpenStudy (mathmale):

.... where \[(x _{1},y _{1,}z _{1})\] ... is the given point, the point that lies in the plane whose equation you wish to find.

OpenStudy (hexiflexidexi):

@mathmale so, I cross the product between the direction vector (how do I find that?) and the vector \[<x-x _{1},y-y _{1},z-z _{1}>, \] (Would any x,y, and z be already okay? Or would they have to be specified?) :) Thanks~

OpenStudy (hexiflexidexi):

Wait, is the direction vector the given one? (2x+y-z=1)?

OpenStudy (hexiflexidexi):

So I cross the vectors (x-0, y-2, z+2) and 2x+y-z=1? (Okai I'm confused) The second one is an equation, right? Or do I find a vector on the given plane?

OpenStudy (baru):

if you are given the equation of a plane, do you know how to find a vector normal to the plane?

OpenStudy (baru):

@HexiFlexiDexi

OpenStudy (hexiflexidexi):

@baru Nope.

OpenStudy (baru):

are you familiar withthis: cross product of two vectors gives us a third vector which is perpendicular to the plane containing the original two vectors

OpenStudy (hexiflexidexi):

Yea :)

OpenStudy (baru):

if i give you the co-ordinates of two points A and B do you know how to find a vector that connects A and B?

OpenStudy (hexiflexidexi):

Yes :) (x1-x2, y1-y2, z1-z2)

OpenStudy (baru):

awesome, now we have the plane 2x+y-z=1 if we could some how find three points A,B and C lying on this plane, we can find two vectors AB and AC lying on the plane. followed so far?

OpenStudy (hexiflexidexi):

Yep!

OpenStudy (baru):

any ideas on how to find three random points on the plane 2x+y-z=1 ?

OpenStudy (baru):

hint: just substitute random easy numbers for any two variables and solve for the third i'll do one, let x=0 and y=0 then 0 +0-z=1 z=-1 so (0,0,-1) is a point on the plane

OpenStudy (baru):

can you find two more?

OpenStudy (hexiflexidexi):

Yea~ (0.5,0,0) and (0,1,0) :)

OpenStudy (baru):

awesome! lets call those points A, B and C can you find the two vectors AB and AC ?

OpenStudy (hexiflexidexi):

AB: (-0.5,1,0) AC: (-0.5,0,-1)?

OpenStudy (baru):

yep that looks right now we have two vectors lying on the given plane, so their cross product should give us a normal vector to the plane. are u following? can you do that?

OpenStudy (hexiflexidexi):

Yep :) (I cross the two vectors?)

OpenStudy (baru):

yep

OpenStudy (hexiflexidexi):

Kay :) (Lemme quickly do this on paper)

OpenStudy (hexiflexidexi):

like this?

OpenStudy (baru):

its right... but you have made a sign mistake, the half in the bottom right corner is negative,

OpenStudy (baru):

so if you correct that, you get N=<-1, -0.5 , 0.5>

OpenStudy (hexiflexidexi):

Ah, right! (Fixed it, to lazy to upload again)

OpenStudy (baru):

now this next part is a bit tricky, two planes are parallel if the have the same normal vector. and a vector is normal(N) to a plane, if ALL vectors lying on the plane are perpendicular to N

OpenStudy (baru):

so let the given point be P=(0 , 2 ,-2) and let X be an unknown point on the required plane X=( x , y , z)

OpenStudy (baru):

can you find the vector PX ? (it will have variables in them)

OpenStudy (hexiflexidexi):

(x,y-2,x+2)?

OpenStudy (baru):

yep... (the x in the end should be z, typo? ) now the vector PX is vector lying on the plane we are trying to find. and we know that this plane has the same normal vector N we just found. so PX is perpendicular to N we know that the dot product of two perpendicular vectors is zero. so can you do the dot product of PX and N ?

OpenStudy (hexiflexidexi):

Yea, typo XD (lemme do it on "paper" again)...

OpenStudy (hexiflexidexi):

Wait, is n the thing we just found using the cross product?

OpenStudy (baru):

yep

OpenStudy (hexiflexidexi):

Like this (sorry for terrible handwriting XD)

OpenStudy (baru):

thats right, now equate that expression to zero (because dot product of two perpendicular vectors is zero) rearrange it and what you get is the equation of the required plane :)

OpenStudy (hexiflexidexi):

:)

OpenStudy (baru):

awesome!! :) now observe closely and compare the answer you found and the given equation. theres a neat shortcut there :)

OpenStudy (hexiflexidexi):

Only the constant part changes? :)

OpenStudy (hexiflexidexi):

@baru :)

OpenStudy (baru):

thats right :)

OpenStudy (baru):

@HexiFlexiDexi

OpenStudy (baru):

2x+y-z=1 is the given equation N=<-1, -0.5, 0.5 > this is the normal vector you found by crossing two vectors on the plane

OpenStudy (baru):

we know that if we multiply a vector with a scalar, it orientation remains the same. so multiply N by "-2" N=<2 , 1 , -1 >